# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
The spectral norm of the residual matrix defined by $\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\bm{\Lambda}^{(i)})$ for $k=0,\dots,p$ is bounded by $\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq f_{k}\lVert\textbf{R}^{(i)}\rVert\,$ where $f_{k}$ is a finite constant and $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$
Proof.
We define the polynomial $P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})$ of degree $k-1$ obtained by the following factorization
$C_{k}(\textbf{H})-C_{k}(\epsilon^{(i)}_{j})=P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})(\textbf{H}-\epsilon^{(i)}_{j}\textbf{I})$ (22)
Consequently the $j$-th column of $\textbf{R}_{k}^{(i)}$ can then be written as
$\left(C_{k}(\textbf{H})-C_{k}(\epsilon^{(i)}_{j})\right)\textbf{x}_{j}=P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})(\textbf{H}-\epsilon^{(i)}_{j}\textbf{I})\textbf{x}_{j}=P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})(\textbf{H}\textbf{x}_{j}-\epsilon^{(i)}_{j}\textbf{x}_{j})$ (23)
where $\textbf{x}_{j}$ is the $j$-th column of $\textbf{X}^{(i)}$ and $\epsilon^{(i)}_{j}$ is the $j$-th diagonal element of $\bm{\Lambda}^{(i)}$. Using the fact that spectral norm of a matrix is bounded by the matrix’s Frobenius norm, we can write the following inequality using Eq. 23
$\lVert\textbf{R}_{k}^{(i)}\rVert\,^{2}\leq\sum_{j=1}^{n}\lVert P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})(\textbf{H}\textbf{x}_{j}-\epsilon^{(i)}_{j}\textbf{x}_{j})\rVert\,^{2}$ (24)
Using the sub-multiplicative property of spectral norm, we have the following
$\lVert P_{k}(\textbf{H},\epsilon^{(i)}_{j})(\textbf{H}\textbf{x}_{j}-\epsilon^{(i)}_{j}\textbf{x}_{j})\rVert\,\leq\lVert P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})\rVert\,\lVert\textbf{H}\textbf{x}_{j}-\epsilon^{(i)}_{j}\textbf{x}_{j}\rVert\,$ (25)
and consequently, we can write
$\lVert\textbf{R}_{k}^{(i)}\rVert\,^{2}$ $\leq\left(\max_{j}\lVert P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})\rVert\,\right)^{2}\sum_{j=1}^{n}\lVert\textbf{H}\textbf{x}_{j}-\epsilon^{(i)}_{j}\textbf{x}_{j}\rVert\,^{2}$ (26)
$\implies\lVert\textbf{R}_{k}^{(i)}\rVert\,$ $\leq\left(\sqrt{n}\max_{j}\lVert P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})\rVert\,\right)\lVert\textbf{R}^{(i)}\rVert\,$ (27)
where in the last step of the above inequality we have used the fact that Frobenius norm of a matrix is bounded by square root of the size of the matrix times the matrix’s spectral norm. Now, defining $f_{k}=\sqrt{n}\max_{j}\lVert P_{k-1}(\textbf{H},\epsilon^{(i)}_{j})\rVert\,$, we have $\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq f_{k}\lVert\textbf{R}^{(i)}\rVert\,$. ∎
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 20 satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\tilde{\eta}_{k}$ is some finite constant that depends on $k$.
Proof.
We argue by induction that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ where $\tilde{\eta}_{k}$ is some non-zero finite constant that depends on $k$. To this end, assuming the induction hypothesis for $k=r-1$ and $k=r$, we have $\lVert\bm{\Delta}_{r-1}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{r-1}\lVert\textbf{R}^{(i)}\rVert\,$ and $\lVert\bm{\Delta}_{r}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,$. Subsequently, the following inequality is true for $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ from Eq. 21
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert a_{r}\right\rvert\,\lVert{\textbf{H}\otimes\underline{\textbf{R}}_{r}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{r}^{(i)}\rVert\,$ (28)
$\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{R}}_{r}^{(i)}\rVert\,\gamma_{m}$ (29)
where we have used Lemma 3.1. We further note that using Lemma 3.2 we can write $\lVert\underline{\textbf{R}}_{r}^{(i)}\rVert\,=\lVert\bm{\Delta}_{r}^{(i)}+\textbf{R}_{r}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,+\lVert\textbf{R}_{r}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,+f_{r}\lVert\textbf{R}^{(i)}\rVert\,$ and consequently we can write
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\Biggl{[}\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\gamma_{m}\tilde{\eta}_{r}+f_{r})\Biggr{]}\gamma_{m}\lVert\textbf{R}^{(i)}\rVert\,$ (30)
Defining $\tilde{\eta}_{r+1}=\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\gamma_{m}\tilde{\eta}_{r}+f_{r})$ we have $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{r+1}\lVert\textbf{R}^{(i)}\rVert\,$. Hence using induction, we argue that $\lVert\bm{\Delta}_{k+1}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k+1}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$
@@
The norm of the residual $\textbf{R}^{(i)}$ is bounded by
$\lVert\textbf{R}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (31)
Proof.
We have, $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$ and $\bm{\Lambda}^{(i)}$ can be written as $\bm{\Lambda}^{(i)}={\textbf{X}^{(i)}}^{\dagger}\textbf{H}\textbf{X}^{(i)}$ allowing us to write $\textbf{R}^{(i)}=(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}$. Let $\hat{\textbf{U}}=\begin{bmatrix}\hat{\textbf{u}}_{1}&\hat{\textbf{u}}_{2}&\dots&\hat{\textbf{u}}_{m}\end{bmatrix}$ be an $m\times m$ unitary matrix with the eigenvectors of H as its columns and $\bm{\Lambda}=\hat{\textbf{U}}^{\dagger}\hat{\textbf{H}}\hat{\textbf{U}}$ is the diagonal matrix of the eigenvalues. Consider a partitioning of $\hat{\textbf{U}}$ as $\begin{bmatrix}\hat{\textbf{U}}_{1}&\hat{\textbf{U}}_{2}\end{bmatrix}$ where $\hat{\textbf{U}}_{1}$ is the $m\times n$ matrix corresponding to the wanted eigenvectors and $\hat{\textbf{U}}_{2}$ is the $m\times(m-n)$ matrix corresponding to the unwanted eigenvectors. Further, considering the partition of $\bm{\Lambda}$ as $\bm{\Lambda}=\begin{bmatrix}\bm{\Lambda}_{1}&0\\ 0&\bm{\Lambda}_{2}\end{bmatrix}$ where $\bm{\Lambda}_{1}$ is the $n\times n$ diagonal matrix comprising of the wanted eigenvalues and $\bm{\Lambda}_{2}$ is the $(m-n)\times(m-n)$ diagonal matrix comprising of the unwanted eigenvalues. We have $\textbf{H}=\hat{\textbf{U}}_{1}\bm{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\bm{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger}$. Consequently we can write
$\lVert\textbf{R}^{(i)}\rVert\,$ $=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}\rVert\,$ (32)
$=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})(\hat{\textbf{U}}_{1}\bm{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\bm{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger})\textbf{X}^{(i)}\rVert\,$ (33)
$\leq\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,\lVert\bm{\Lambda}_{1}\rVert\,\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,+\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{2}\rVert\,\lVert\bm{\Lambda}_{2}\rVert\,\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,$ (34)
@@
$\lVert\textbf{R}^{(i)}\rVert\,$ $\leq\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}(\lVert\bm{\Lambda}_{1}\rVert\,+\lVert\bm{\Lambda}_{2}\rVert\,)\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (35)
@@
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (36)
If this inequality is satisfied for $i=i_{0}$ then it holds for all $i>i_{0}$
Proof.
As a consequence of Lemmas 3.2, 3.3 and 3.4 we have $\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from Eq. 4 we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (37)

### v3

@@ 3.2.2 Proposed residual-based Chebyshev filtering approach
The spectral norm of the residual, $\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)})$ for $k=0,\dots,p$, is bounded by $\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq f_{k}\lVert\textbf{R}^{(i)}\rVert\,$ where $f_{k}$ is a finite constant and $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$
Proof 3.12.
Let $C_{k}(x)=\sum_{j=0}^{k}\alpha_{j}x^{j}$, we can now write
$\textbf{R}_{k}^{(i)}$ $=\sum_{j=1}^{k}\alpha_{j}\left(\textbf{H}^{j}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}^{j}\right)=\sum_{j=1}^{k}\alpha_{j}\sum_{r=0}^{j-1}\textbf{H}^{j-r-1}\left(\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\right){\boldsymbol{\Lambda}^{(i)}}^{r}$
Using submultiplicativity of the spectral norm we can now write
$\lVert\textbf{R}_{k}^{(i)}\rVert\,$ $\leq\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,=f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
where we have defined $f_{k}=\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,$.
@@
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using lower precision arithmetic in matrix-products of the recurrence relation Eq. 16 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\tilde{\eta}_{k}$ is some finite constant that depends on $k$.
Proof 3.14.
Using Eq. 14 and Eq. 16 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(17) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}\right)$
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=\boldsymbol{\Delta}_{1}^{(i)}=0$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with $\textbf{e}_{k}^{(i)}={\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}$. From Lemma 3.3 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{H}\otimes\underline{\textbf{R}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{R}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\underline{\textbf{R}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{R}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{R}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\gamma_{m}f_{k}\lVert\textbf{H}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$. Using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=0$, $h_{0}=\gamma_{m}\lVert\textbf{H}\rVert\,$ and $h_{1}^{(i)}=\gamma_{m}f_{k}\lVert\textbf{H}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ with $\tilde{\eta}_{k}=f_{k}\lVert\textbf{H}\rVert\,g_{k-1}^{1}$.
@@
The norm of the residual $\textbf{R}^{(i)}$ satisfies $\lVert\textbf{R}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
Proof 3.16.
We have, $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$ and the subspace diagonalization performed in the Rayleigh-Ritz projection step associated with $(i-1)^{th}$ iteration allows one to write $\boldsymbol{\Lambda}^{(i)}$ as $\boldsymbol{\Lambda}^{(i)}={\textbf{X}^{(i)}}^{\dagger}\textbf{H}\textbf{X}^{(i)}$. Subsequently we can write the residual $\textbf{R}^{(i)}=(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}$. We note that $\textbf{H}=\hat{\textbf{U}}_{1}\boldsymbol{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\boldsymbol{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger}$ and consequently we can write
$\lVert\textbf{R}^{(i)}\rVert\,=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}\rVert\,=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})(\hat{\textbf{U}}_{1}\boldsymbol{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\boldsymbol{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger})\textbf{X}^{(i)}\rVert\,\\ \leq\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,\lVert\boldsymbol{\Lambda}_{1}\rVert\,\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,+\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{2}\rVert\,\lVert\boldsymbol{\Lambda}_{2}\rVert\,\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,$
@@
(18) $\lVert\textbf{R}^{(i)}\rVert\,$ $\leq\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}(\lVert\boldsymbol{\Lambda}_{1}\rVert\,+\lVert\boldsymbol{\Lambda}_{2}\rVert\,)\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
@@
(19) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
and if this inequality is satisfied for $i=i_{0}$ then it holds for all $i>i_{0}$
Proof 3.18.
From Lemmas 3.11, 3.13, and 3.15 we have $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from Eq. 3 we require
(20) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$